# Progressive-Proximity Bit-Flipping for Decoding Surface Codes

### v1

@@ Abstract
Topological quantum codes, such as toric and surface codes, are excellent candidates for hardware implementation due to their robustness against errors and their local interactions between qubits. However, decoding these codes efficiently remains a challenge: existing decoders often fall short of meeting requirements such as having low computational complexity (ideally linear in the code’s blocklength), low decoding latency, and low power consumption. In this paper we propose a novel bit-flipping (BF) decoder tailored for toric and surface codes. We introduce the proximity vector as a heuristic metric for flipping bits, and we develop a new subroutine for correcting a particular class of harmful degenerate errors. Our algorithm achieves linear complexity growth and it can be efficiently implemented as it only involves simple operations such as bit-wise additions, quasi-cyclic permutations and vector-matrix multiplications. The proposed decoder shows a decoding threshold of $7.5\%$ for the 2D toric code and $7\%$ for the rotated planar code over the binary symmetric channel.
@@ I Introduction
Bit flipping (BF) decoders are a class of iterative decoding algorithms known to be very fast and efficient [12], although generally they provide lower performance than BP. In general, BF decoders do not perform well on surface codes, mainly because of the very low column weight of the parity check matrix, and also because of error degeneracy, which refers to the property of quantum codes where multiple error patterns of the same weight can correspond to the same syndrome. Nevertheless, because of its extremely low complexity, BF is still attractive in the scenario of decoding topological codes: indeed, the latency constraint that a quantum decoder should meet is very tight, to prevent the so-called backlog problem [13]. Moreover, it is essential that the decoder has a low power consumption, as it has to be embedded in a cryogenic environment with a strict power budget [14].
@@
In this paper, we develop a BF algorithm which is capable of decoding surface codes. In the proposed approach, instead of considering the number of unsatisfied checks as in conventional BF, each bit is assigned a heuristic weight which is the entry of what we call proximity vector. The proximity vector is the sum of different contributions called individual influences, and each individual influence is associated with an unsatisfied check. Ultimately, bits connected with checks with low entries in the proximity vector will be flipped first, while bits connected with checks with high entries in the proximity vector will be flipped last. After each flip, the proximity vector is updated in an efficient manner. To deal with high-weight consecutive errors, i.e., errors occurring on a set of adjacent qubits, which are particularly harmful for iterative decoders, we design an iterative matching procedure that runs after BF and that is able to correct these errors. We show that our decoder has an asymptotic complexity of $\mathcal{O}(n)$, where $n$ is the code’s blocklength, and we provide simulation results that show a comparison, in terms of performance, with MWPM, UF and traditional BF, as well as the average number of decoding iterations as a function of the blocklength.
@@
The rest of the paper is organized as follows. In Section II we introduce the preliminaries of quantum error correction. In Section III we present an overview of the main decoding algorithms for topological codes. In Section IV we define the proximity vector and how it can be computed efficiently. Section V provides a detailed description of our proposed decoder. In Section VI we carry out a complexity analysis and a comparison with other state-of-art decoders. Finally, Section VII presents simulation results.
@@ II-A Stabilizer formalism
A stabilizer group $\mathcal{S}$ is an Abelian subgroup of $\mathcal{G}_{n}$, and an $\llbracket n,k,d\rrbracket$ stabilizer code is a $2^{k}$-dimensional subspace $\mathcal{C}$ of the Hilbert space $(\mathbb{C}^{2})^{\otimes n}$ that satisfies the condition $S_{i}\ket{\Psi}=\ket{\Psi},\ \forall\ S_{i}\in\mathcal{S},\ket{\Psi}\in\mathcal{C}$. Thus, the code $\mathcal{C}$ is defined as the common $+1$-eigenspace of the stabilizer group $\mathcal{S}$. The stabilizer group $\mathcal{S}$ is generated by a set of $n-k$ independent generators $S_{1},...,S_{n-k}$ that can be represented using a matrix $\mathbf{S}$, called the stabilizer matrix, whose $(i,j)$ element is given by the Pauli operator corresponding to the $j$-th qubit in the $i$-th stabilizer. The minimum distance $d$ is defined as the minimum weight of an element of $\mathcal{S}$.
Applying the mapping (1) to the stabilizer matrix $\mathbf{S}$, we obtain a $(n-k)\times 2n$ binary matrix:
@@ II-B Error model and syndrome
Obviously, if $\mathbf{e}\in\mathcal{S}$ we have $\mathbf{s}=\mathbf{0}$. More generally, any Pauli error can be expressed as a combination of a true error and a stabilizer such that $\mathbf{e}=\mathbf{e}_{t}+\mathbf{S}_{i}$, with $\mathbf{S}_{i}\in\mathcal{S}$; since, by definition, elements of the stabilizer group $\mathcal{S}$ commute with each other (thus, their syndrome is zero), the syndrome $\mathbf{s}$ is only dependent on the true error $\mathbf{e}_{t}$; nonetheless, this also means that for every true error $\mathbf{e}_{t}$, there exist a set of Pauli errors $\mathcal{E}=\mathbf{e}_{t}+\mathcal{S}$ that will lead to the same syndrome. This phenomenon is known as error degeneracy. It is well known that, if the minimum distance of a stabilizer code is much higher than the weight of its stabilizer elements, there will be many degenerate errors of the same weight [15].
@@ II-C Calderbank-Shor-Steane codes
An $\llbracket n,k_{X}-k_{Z},d\rrbracket$ Calderbank-Shor-Steane (CSS) code $\mathcal{C}$ is a stabilizer code constructed using two classical $[n,k_{X},d_{X}]$ and $[n,k_{Z},d_{Z}]$ codes $C_{X}$ and $C_{Z}$, respectively, where $d\geq\leavevmode\nobreak\ \mathrm{min}\{d_{X},d_{Z}\}$ and $C_{Z}\subset C_{X}$ [16]. Note that $k_{X}$, $k_{Z}$ and $d_{X}$, $d_{Z}$ correspond to the dimensions and minimum distances of $C_{Z}$ and $C_{X}$, respectively. The stabilizer matrix of the CSS code $\mathcal{C}$ has the form
@@
II-D Toric and surface codes
Toric codes [1] are a widely known class of CSS codes. These are derived from the tessellation of the topological surface of a torus in squares, in such a way as to form a lattice. Generally, a toric code is characterized by a $L\times L$ lattice, where $L$ is the size of the horizontal (or vertical) dimension. To an $L\times L$ lattice corresponds a $\llbracket 2L^{2},2,L\rrbracket$ quantum CSS code. The $\mathbf{H}_{X}$ matrix is the incidence matrix between vertices (check nodes) and edges (variable nodes), and the $\mathbf{H}_{Z}$ matrix is the incidence matrix between squares (check nodes) and edges (variable nodes). The two Tanner graphs corresponding to $\mathbf{H}_{X}$ and $\mathbf{H}_{Z}$ are identical and they obviously correspond to a torus as well, as depicted in Fig. 1.
Surface codes are similar to toric codes, except that the lattice has now open boundaries; in such a way, the code can be implemented on a two-dimensional quantum chip [19]. Although many variants of surface codes have been studied, we focus on the $\llbracket L^{2},1,L\rrbracket$ rotated planar codes [19], which are obtained by a $45\degree$ rotation of an $L\times L$ lattice, after the pruning of unnecessary qubits.
0000111144444444666633333333111122225555555577772222003333333388889999666612121212777710101010131313138888111111111414141466669999151515150000111116161616111122221717171722220015151515
Figure 1: Tanner graph of the $\llbracket 18,2,3\rrbracket$ toric code. Black values correspond to variable node labels, while red values are check node labels.
@@ III-A Minimum weight perfect matching
The decoding of toric and surface codes is usually carried out using the MWPM algorithm [2]. After measuring the syndrome, a complete graph is constructed, where each vertex represents an unsatisfied check; each edge is assigned a weight, which is the minimum number of qubits separating the two checks connected by the edge; this weight assignment is done using Dijkstra’s algorithm [20]. Then, the blossom algorithm [21] is run on this graph, with the goal of finding the minimum weight perfect matching. A matching on a graph is a subset of edges such that any two edges in this subset do not share any common vertex; a perfect matching is a matching such that every vertex in the graph is connected to exactly one edge. A minimum weight perfect matching is a perfect matching such that the sum of all the weights of the edges of the matching is minimum among all the possible perfect matchings. Since the overall complexity of the MWPM decoder on an $L\times L$ lattice is $\mathcal{O}(L^{6}\log L)$ [22], many efficient implementations have been proposed to reduce this complexity: for instance, in [2] a fully parallel implementation is proposed, able to run with a complexity of $\mathcal{O}(1)$, given a uniform 2-D array of finite speed processing elements and an external memory able to store all the detection events and matching data for the entire duration of a hypothetical computation, while Higgott et al. proposed, in [22], that each unsatisfied check could be restricted to be matched within a local neighborhood, or that Dijkstra’s algorithm could be optimized to avoid an all-to-all search [23].
III-B Union-find
The main competitor of MWPM is the UF decoder [3]. This decoding scheme is also divided into two parts: first, the syndrome validation process is carried out, where clusters are defined to initially correspond to each unsatisfied check, and are then iteratively grown to absorb the nearest check neighbors. A cluster stops growing when it contains an even number of unsatisfied checks, and two clusters merge together if they touch each other. After all the clusters have stopped growing, the peeling decoder is run on each cluster; it is possible to treat errors as erasures because the error location is known after the syndrome validation step, given that each cluster is a spanning tree; otherwise, clusters are split in order to obtain spanning trees [24]. The complexity of the UF decoder, in its efficient implementation proposed in [3], is almost linear in the blocklength.
III-C Message-passing decoders
@@
Message passing decoders such as BP are, in principle, not suitable for topological codes, mainly because of the presence of weight-4 symmetric stabilizers that behave as trapping sets for the decoder [25]; because of this, several techniques have been developed to improve the performance of BP on these codes. One of the most common approaches is to apply the well-known ordered statistics technique after several rounds of BP [10]; this decoder is known as BP-OSD and has a complexity of $\mathcal{O}(n^{3})$, thus making it infeasible for large-scale implementation. More recent contributions have shown that a serial scheduling of the messages, along with modifications of the update rules and re-initialization, lead to a significant increase of the performance at the cost of increasing the decoding latency [11, 26].
III-D Motivation for our work
For the majority of decoders proposed for topological codes, low decoding error probability comes with the price of increased decoding complexity. MWPM, for instance, has a complexity that is cubic in the blocklength and it is known to be infeasible in practice; the efficient implementations available, in particular those of [22] and [23], offer a significant speedup that still vanishes for large blocklength. The UF decoder, on the other hand, is able to offer an almost-linear time complexity by exploiting an efficient implementation, and its performance is somewhat close to that of MWPM; nevertheless, parallel implementations of this decoder come with several challenges: in [27], for instance, it is shown that the worst-case complexity rises to $\mathcal{O}(L^{3})$ when the decoder is implemented using parallel computational units that exchange messages with their nearest neighbors. Indeed, besides low asymptotic complexity, it is important that a decoder also has low decoding latency. In this sense, message-passing decoders such as BP are known to be excellent, as the update rules are local and equal for each node type, and can be easily implemented in parallel. Nevertheless, algorithms like BP-OSD fail to achieve linear complexity.
@@ III-D Motivation for our work
In this scenario, it seems reasonable to revisit the BF decoder, as it not only exhibits a complexity linear in $n$, but also utilizes extremely simple update rules that make it much faster than BP. However, because of the low degree of variable nodes in the Tanner graph of the considered quantum codes (toric codes are regular with variable degree of 2, while surface codes also have degree-1 nodes), and because of error degeneracy, we need to make significant modifications to the traditional BF algorithm in order to decode toric and surface codes. We will show that these modifications do not impact on the asymptotic complexity of the decoder, and that they significantly improve the performance of the decoder comparing to the one of the traditional BF; we also compare the performance of our algorithm with the ones of MWPM and UF.
@@ IV The proximity vector
for $\ell=1,2,...,D$. We call $\boldsymbol{\nu}^{(\ell)}(c_{j})$ the proximity influence of $c_{j}$ on qubits (or variable nodes) of depth $\ell$, and $\boldsymbol{\gamma}^{(\ell)}(c_{j})$ the proximity influence of $c_{j}$ on checks of depth $\ell$.
@@
Since the value $D$ will be fixed in the rest of the paper, we simply refer to the influences as $\boldsymbol{\nu}(c_{j})$ and $\boldsymbol{\gamma}(c_{j})$, and to the vectors as $\boldsymbol{\nu}$ and $\boldsymbol{\gamma}$. By construction, the values of the influences $\boldsymbol{\nu}(c_{j})$ and $\boldsymbol{\gamma}(c_{j})$ are highest for the nearest neighbors of $c_{j}$, and decrease with increasing distance of a variable (check) node from $c_{j}$. The proximity vector on qubits $\boldsymbol{\nu}$ and on checks $\boldsymbol{\gamma}$ is the natural superposition of all of the influences exerted by all of the unsatisfied checks. As a result, variable and check nodes which are more distant from unsatisfied nodes (we say more isolated) will have a lower weight assigned, while variable and check nodes that are near to many unsatisfied checks will have higher weights assigned. Note that this definition of proximity vector is heuristic, and alternative definitions can be provided to capture the individual influence of a check node. Our definition of proximity vector is convenient for the reason that it only involves integer values, thus it can be stored using a low fixed (and low) number of bits.

### v2

@@ Abstract
Topological quantum codes, such as toric and surface codes, are excellent candidates for hardware implementation due to their robustness against errors and their local interactions between qubits. However, decoding these codes efficiently remains a challenge: existing decoders often fall short of meeting requirements such as having low computational complexity (ideally linear in the code’s blocklength), low decoding latency, and low power consumption. In this paper we propose a novel bit-flipping (BF) decoder tailored for toric and surface codes. We introduce the proximity vector as a heuristic metric for flipping bits, and we develop a new subroutine for correcting degenerate multiple errors on adjacent qubits. Our algorithm has quadratic complexity growth and it can be efficiently implemented as it does not require operations on dynamic memories, as do state-of-art decoding algorithms such as minimum weight perfect matching or union find. The proposed decoder shows a decoding threshold of $7.5\%$ for the 2D toric code and $7\%$ for the rotated planar code over the binary symmetric channel.
@@ I Introduction
Bit flipping (BF) decoders are a class of iterative decoding algorithms known to be very fast and efficient [12], although generally they provide lower performance than BP. In general, BF decoders do not perform well on surface codes, mainly because of the very low column weight of the parity check matrix, and also because of error degeneracy, which refers to the property of quantum codes where multiple error patterns of the same weight can correspond to the same syndrome. Nevertheless, because of its extremely low complexity, BF is still attractive in the scenario of decoding topological codes; indeed, the latency constraint that a quantum decoder should meet is very tight, to prevent the so-called backlog problem [13]. Moreover, it is essential that the decoder has a low power consumption, as it has to be embedded in a cryogenic environment with a strict power budget [14]. This makes hardware implementation of the decoder a crucial aspect; unfortunately, state-of-art decoders such as MWPM or UF do not allow an efficient hardware implementation, mainly because they make use of complex data structures, which require random memory access and dynamic memory allocation, which are known to be less efficient as they introduce latency and increase circuit complexity as well as power consumption. In contrast, the BF algorithm, due to its simplicity, only requires static memories with fixed access, thus having a distinct advantage for hardware implementation.
@@
In this paper, we develop a BF algorithm which is capable of decoding surface codes. In the proposed approach, instead of considering the number of unsatisfied checks as in conventional BF, each bit is assigned a heuristic weight which is the entry of what we call proximity vector. The proximity vector is the sum of different contributions called individual influences, and each individual influence is associated with an unsatisfied check. Ultimately, bits connected with checks with low entries in the proximity vector will be flipped first, while bits connected with checks with high entries in the proximity vector will be flipped last. After each flip, the proximity vector is updated in an efficient manner. To deal with high-weight consecutive errors, i.e., errors occurring on a set of adjacent qubits, which are particularly harmful for iterative decoders, we design an iterative matching procedure that runs after BF and that is able to correct these errors. We show that our decoder has an asymptotic complexity of $\mathcal{O}(n^{2})$, where $n$ is the code’s blocklength, and we provide simulation results that show a comparison, in terms of performance, with MWPM, UF and traditional BF. Although our decoder is not able to correct up to the minimum distance of the code, we show that it has a good performance while also being more feasible for hardware implementation than UF and MWPM.
@@
The rest of the paper is organized as follows. In Section II we introduce the preliminaries of quantum error correction. In Section III we present an overview of the main decoding algorithms for topological codes. In Section IV we define the proximity vector and how it can be computed efficiently. Section V provides a detailed description of our proposed decoder. In Section VI we carry out a complexity analysis and a comparison with other state-of-art decoders. In Section VII we analyze the hardware implementation aspects for UF, MWPM and the proposed decoder. Finally, Section VIII presents simulation results.
@@ II-A Stabilizer formalism
A stabilizer group $\mathcal{S}$ is an Abelian subgroup of $\mathcal{G}_{n}$, and an $\llbracket n,k,d\rrbracket$ stabilizer code is a $2^{k}$-dimensional subspace $\mathcal{C}$ of the Hilbert space $(\mathbb{C}^{2})^{\otimes n}$ that satisfies the condition $S_{i}\ket{\Psi}=\ket{\Psi},\ \forall\ S_{i}\in\mathcal{S},\ket{\Psi}\in\mathcal{C}$. Thus, the code $\mathcal{C}$ is defined as the common $+1$-eigenspace of the stabilizer group $\mathcal{S}$. The stabilizer group $\mathcal{S}$ is generated by a set of $n-k$ independent generators $S_{1},...,S_{n-k}$ that can be represented using a matrix $\mathbf{S}$, called the stabilizer matrix, whose $(i,j)$ element is given by the Pauli operator corresponding to the $j$-th qubit in the $i$-th stabilizer. The minimum distance $d$ is defined as the minimum weight of an element of $N(\mathcal{S})\setminus\mathcal{S}$, where $N(\mathcal{S})$ is the normalizer [^1] The normalizer subgroup $N(\mathcal{S})$ of a group $\mathcal{S}$ is the subgroup of $\mathcal{S}$ which is invariant under conjugation. of $\mathcal{S}$. Elements in $N(\mathcal{S})\setminus\mathcal{S}$ are also called logical operators.
Applying the mapping (1) to the stabilizer matrix $\mathbf{S}$, we obtain an $(n-k)\times 2n$ binary matrix:
@@ II-B Error model and syndrome
Obviously, if $\mathbf{e}\in\mathcal{S}$ we have $\mathbf{s}=\mathbf{0}$. More generally, any Pauli error can be expressed as a combination of a true error $\mathbf{e}_{t}$, a stabilizer $\mathbf{S}_{i}$ and a logical operator $\mathbf{L}_{j}$ such that $\mathbf{e}=\mathbf{e}_{t}+\mathbf{S}_{i}+\mathbf{L}_{j}$, with $\mathbf{S}_{i}\in\mathcal{S}$ and $\mathbf{L}_{j}\in N(\mathcal{S})\setminus\mathcal{S}$; since, by definition, elements of $\mathbf{L}_{j}$ and $\mathcal{S}_{i}$ commute with each other (thus, their syndrome is zero), the syndrome $\mathbf{s}$ is only dependent on the true error $\mathbf{e}_{t}$; moreover, any error estimate of the type $\hat{\mathbf{e}}=\mathbf{e}_{t}+\mathbf{S}_{i}$ is a valid error estimate. This phenomenon is known as error degeneracy. It is well known that, if the minimum distance of a stabilizer code is much higher than the weight of its stabilizer elements, there will be many degenerate errors of the same weight [15].
@@ II-C Calderbank-Shor-Steane codes
An $\llbracket n,k_{X}-k_{Z},d\rrbracket$ Calderbank-Shor-Steane (CSS) code $\mathcal{C}$ is a stabilizer code constructed using two classical $[n,k_{X},d_{X}]$ and $[n,k_{Z},d_{Z}]$ codes $C_{X}$ and $C_{Z}$, respectively, where $d\geq\leavevmode\nobreak\ \mathrm{min}\{d_{X},d_{Z}\}$ and $C_{Z}\subset C_{X}$ [16]. Note that $k_{X}$, $k_{Z}$ and $d_{X}$, $d_{Z}$ correspond to the dimensions and minimum distances of $C_{Z}$ and $C_{X}$, respectively. The parity check matrix of the CSS code $\mathcal{C}$ has the form
@@
II-D Surface and toric codes
Surface codes [1] are a widely known class of CSS codes. These are derived from the tessellation of the topological surface in squares, in such a way as to form a lattice. Generally, a surface code is characterized by an $L\times L$ lattice, where $L$ is the size of the horizontal (or vertical) dimension. In the lattice we can identify vertices as $X$ checks, edges as qubits and squares as $Z$ checks. Although there are several types of surface codes, in this paper we focus on the rotated planar codes [19], with parameters $\llbracket L^{2},1,L\rrbracket$ ($L$ being the size of the lattice) for the description of the decoder. The matrix $\mathbf{H}_{X}$ is the incidence matrix between vertices and edges, and the matrix $\mathbf{H}_{Z}$ is the incidence matrix between squares and edges. The two Tanner graphs corresponding to $\mathbf{H}_{X}$ and $\mathbf{H}_{Z}$ are isomorphic and they obviously correspond to a lattice as well, as depicted in Fig. 1 for the rotated planar code. We also consider toric codes, which are again characterized by a $L\times L$ lattice, such that the edges and vertices at the boundaries coincide. The toric code has parameters $\llbracket 2L^{2},2,L\rrbracket$.
000011112222111133334444555522226666777788883333
Figure 1: Tanner graph for the $\llbracket 9,1,3\rrbracket$ rotated planar code. Circles (with black labels) correspond to variable nodes, while squares (with red labels) correspond to check nodes.
@@ III Decoding of toric and surface codes
The decoding of toric and surface codes is usually carried out using the MWPM algorithm [2]. After measuring the syndrome, a complete graph is constructed, where each vertex represents an unsatisfied check; each edge is assigned a weight, which is the minimum number of qubits separating the two checks connected by the edge; this weight assignment is done using Dijkstra’s algorithm [20]. Then, the blossom algorithm [21] is run on this graph, with the goal of finding the minimum weight perfect matching. Since the overall complexity of the MWPM decoder on an $L\times L$ lattice is $\mathcal{O}(L^{6}\log L)$ [22], many efficient implementations have been proposed to reduce this complexity: for instance, in [2] a fully parallel implementation is proposed, able to run with a complexity of $\mathcal{O}(1)$, given a uniform 2-D array of finite speed processing elements and an external memory able to store all the detection events and matching data for the entire duration of a hypothetical computation, while Higgott et al. proposed, in [22], that each unsatisfied check could be restricted to be matched within a local neighborhood, or that Dijkstra’s algorithm could be optimized to avoid an all-to-all search [23]. The main competitor of MWPM is the UF decoder [3]. This decoding scheme is also divided into two parts: first, the syndrome validation process is carried out, where clusters are defined to initially correspond to each unsatisfied check, and are then iteratively grown to absorb the nearest check neighbors. A cluster stops growing when it contains an even number of unsatisfied checks, and two clusters merge together if they touch each other. After all the clusters have stopped growing, the peeling decoder is run on each cluster. The complexity of the UF decoder, in its efficient implementation proposed in [3], is almost linear in the blocklength. There are also several alternative approaches to UF, which are able to improve the decoding latency. For instance, in [24], the authors propose a spanning tree (ST) matching decoder which is able to correct up to the code distance. In addition, they also propose a lower complexity decoder, called the rapid-fire decoder, which is able to improve the decoding latency at the cost of some performance degradation.
@@
Message passing decoders such as BP are, in principle, not suitable for topological codes, mainly because of the presence of weight-4 symmetric stabilizers that behave as trapping sets for the decoder [25]; because of this, several techniques have been developed to improve the performance of BP on these codes. One of the most common approaches is to apply the well-known ordered statistics technique after several rounds of BP [10]; this decoder is known as BP-OSD and has a complexity of $\mathcal{O}(n^{3})$, thus making it infeasible for large-scale implementation.
For the majority of decoders proposed for topological codes, low decoding error probability comes with the price of increased decoding complexity. MWPM, for instance, has a complexity that is cubic in the blocklength and it is known to be infeasible in practice; the efficient implementations available, in particular those of [22] and [23], offer a significant speedup that still vanishes for large blocklength. Another major issue in decoders for topological codes, which is often overlooked, is the hardware implementation aspect of the algorithms. In fact, the vast majority of these decoders turn out to be inefficient when implemented on FPGAs. It is well-known that fixed and predictable access to pre-allocated memory is much faster than random access to irregular, pointer-based dynamic memory; in particular, FPGAs are particularly suitable for the former type of memory access. Moreover, modifications of dynamic structures (like adding nodes to a tree, or modifying its connectivity) are also operations which are known to introduce a high amount of latency, even when parallelized, due to the high overhead required. Another aspect to consider is that the usage of dynamic memories is usually more expensive in terms of energy consumption, as they often require additional overhead for their management, as well as more complex hardware. Also, energy consumption is a critical aspect of any quantum error correction decoder, as it runs on a chip which is embedded in a cryogenic environment with a strict power budget. MWPM and spanning tree decoders, for example, make use of Dijkstra’s algorithm to construct syndrome graphs on which they perform the perfect matching. Dijkstra’s algorithm is not desirable for an efficient FPGA implementation, primarily due to its inherently sequential nature and dependence on dynamic data structures (priority queues).
Similarly, the UF decoder relies on dynamic, pointer-based data structures and complex tree manipulations. The two operations, $\mathtt{union()}$ and $\mathtt{find()}$, often involve traversing and modifying tree structures, which require frequent pointer updates and non-uniform memory access patterns.
@@
In this scenario, it seems reasonable to revisit the BF decoder which utilizes extremely simple update rules that make it much easier to implement in hardware such as FPGA. However, because of the low degree of variable nodes in the Tanner graph of the considered quantum codes (toric codes are regular with variable degree of 2, while surface codes also have degree-1 nodes), and because of error degeneracy, we need to make significant modifications to the traditional BF algorithm in order to decode toric and surface codes. We will show that, although these modifications impact on the asymptotic complexity of the decoder, they significantly improve the performance of the decoder comparing to that which uses traditional BF; at the same time, our algorithm does not make use of any dynamic data structures, and only utilizes elementary operations on arrays of fixed length which can be pre-allocated in memory, thus making our decoder particularly suitable for efficient hardware implementation compared to the alternatives. Nevertheless, the goal of this paper is to merely present and explain the algorithm, without presenting an explicit hardware implementation; an optimized hardware implementation, together with a detailed comparison with those of other decoders, requires a specific in-depth study and is therefore left for future work.
@@ IV The proximity vector
for $\ell=1,2,...,D$, $D$ being a fixed positive integer. We call $\boldsymbol{\nu}^{(\ell)}(c_{j})$ the proximity influence of $c_{j}$ on qubits (or variable nodes) of depth $\ell$, and $\boldsymbol{\gamma}^{(\ell)}(c_{j})$ the proximity influence of $c_{j}$ on checks of depth $\ell$.
@@
Since the value of $D$ will be fixed in the rest of the paper, we simply refer to the proximity influences as $\boldsymbol{\nu}(c_{j})$ and $\boldsymbol{\gamma}(c_{j})$, and to the vectors as $\boldsymbol{\nu}$ and $\boldsymbol{\gamma}$. By construction, the values of the influences $\boldsymbol{\nu}(c_{j})$ and $\boldsymbol{\gamma}(c_{j})$ are highest for the nearest neighbors of $c_{j}$, and decrease with increasing distance of a variable (check) node from $c_{j}$. To see this, it is convenient to write $\boldsymbol{\nu}^{(D)},\boldsymbol{\gamma}^{(D)}$ in the following recursive way. Let $\mathcal{N}(c_{j})\subseteq V$ be the neighborhood of check node $c_{j}$, and $\mathcal{N}(v_{j})\subseteq C$ the neighborhood of variable node $v_{j}$; we define $\mathcal{N}^{d}(c_{j})\subseteq V$ to be the depth-$d$ neighborhood of $c_{j}$, i.e., the shortest path between $c_{j}$ and its depth-$d$ neighbors has length $12d-$ 1; we can also define $\mathcal{N}^{d}(v_{j})\subseteq C$ in a similar way. For every check node $c_{k}\in C$ and variable node $v_{q}\in V$:
$\begin{cases}\gamma_{k}^{(i)}(c_{j})=\gamma_{k}^{(i-1)}(c_{j})+\sum_{v_{z}\in\mathcal{N}(c_{k})}\nu_{z}^{(i-1)}(c_{j})\\ \nu_{q}^{(i)}(c_{j})=\nu_{q}^{(i-1)}(c_{j})+\sum_{c_{z}\in\mathcal{N}(v_{q})}\gamma_{z}^{(i-1)}(c_{j}).\end{cases}$ (12)
For $i=0$ we have $\nu_{q}^{(0)}(c_{j})=1$ for all variable nodes $v_{q}\in\mathcal{N}(c_{j})$ and 0 otherwise; for $i=1$ we have $\nu_{q}^{(1)}(c_{j})=1$ for all variable nodes $v_{q}\in\mathcal{N}^{2}(c_{j})$; moreover, $\nu_{q}^{(1)}(c_{j})>1$ for all variable nodes $v_{q}\in\mathcal{N}(c_{j})$. Let $r>i+1$ and $m<i+1$; in general we have $\nu_{q}^{(i)}(c_{j})=0$ for all variable nodes in $\mathcal{N}^{r}(c_{j})$, $\nu_{q}^{(i)}(c_{j})=1$ for all variable nodes in $\mathcal{N}^{i+1}(c_{j})$, and $\nu_{q}^{(i)}(c_{j})>1$ for variable nodes in $\mathcal{N}^{m}(c_{j})$. In other words, the value of the proximity influence for each node at distance $12d-$ 1 from $c_{j}$ remains 00 until the $d+1$ iteration, and then increases according to (12). This means that, fixing the number of iterations, the closer is a node to $c_{j}$, the higher is the proximity influence for that node. A similar argument can be made for $\boldsymbol{\gamma}(c_{j})$.
The proximity vector on qubits $\boldsymbol{\nu}$ and on checks $\boldsymbol{\gamma}$ is the natural superposition of all of the influences exerted by all of the unsatisfied checks. As a result, variable and check nodes which are more distant from unsatisfied nodes (we say more isolated) will have a lower weight assigned, while variable and check nodes that are near to many unsatisfied checks will have higher weights assigned. Note that this definition of proximity vector is heuristic, and alternative definitions can be provided to capture the individual influence of a check node. Our definition of proximity vector is convenient for the reason that it only involves integer values, thus it can be stored using a fixed (and low) number of bits.